Gemini South telescope has resumed nighttime operations, after having remained closed due to bad winter weather. Only IGRINS and Zorro are available, and the scheduled telescope shutdown will start next Monday, August 01.
Science Operations Announcements
Gemini South telescope has remained closed the entire past week after heavy snowfall over Pachon. The large quantity of snow accumulated over the summit during the two events between July 9 and July 15 prevented the telescope from being open, with large accumulation of snow and ice on the roads as well. Additionally, commercial power was lost and its return is expected sometime during this week, although there is still no certainty about that.
Gemini South Telescope will remain closed through at least July 25th due to a large accumulation of snow and ice on the dome and roads. It is uncertain when the roads will be passable and a crew can travel to the summit to begin clearing the large amount of accumulated snow on the dome. Additionally, commercial power was lost and its return is also not expected until after the roads are cleared.
A patch release of DRAGONS, v3.0.3, is now available. DRAGOSN v3.0.3 is a patch release that offers an automated handling of the GMOS-S amplifier #5 that failed in January. The patch also includes a few other bug fixes.
The 2022B Gemini Observing Tool is now available for download. This update is required to access the Gemini Observing Databases.
See the the Release Notes for a complete list of changes.
The Gemini observing databases will be offline for several hours on Wednesday, June 1 while we perform software updates.
During this time the following services will be unavailable:
- Observing Tool synchronizations and file transfers
- Phase I Tool guide star probability queries
- Target of Opportunity triggers
The new 2022B Observing Tool will be available after the update and will be required to access the observing databases.
A patch release for Gemini IRAF v1.14 is available.
The replacement of two filters and the reshuffling of another in Flamingos 2 necessitated additions and modifications to lookup tables and modifications to the code. This patch release contains only the affected files. The patch is to be applied on-top of an already installed Gemini IRAF v1.14 package.
Installation instructions and content of the patch are available in gemini_v114_patch1.txt.
Today Thursday March 10th we will reinstall the repaired P2WFS in the telescope. This requires removal and reinstall of the A&G which takes more than a day and thus tonight March 10th the telescope will be in shutdown and we expect that tomorrow night we will be back in operation.
It has been found after extensive troubleshooting that the Peripheral Wave Front Sensor 2 (P2WFS) has a mechanical issue and is out of order. To address this issue the P2WFS will have to be replaced in the telescope by a dummy unit while it is being repaired. We are planning for this intervention starting Monday February 21st and the removal and installation of the dummy unit will take two days and thus the telescope will not be available from the Night February 21st to 22nd.